# spacing=nonfrench

### v1

@@ Abstract
Pose-driven human image animation has achieved tremendous progress, enabling the generation of vivid and realistic human videos from just one single photo. However, it conversely exacerbates the risk of image misuse, as attackers may use one available image to create videos involving politics, violence and other illegal content. To counter this threat, we propose Dormant, a novel protection approach tailored to defend against pose-driven human image animation techniques. Dormant applies protective perturbation to one human image, preserving the visual similarity to the original but resulting in poor-quality video generation. The protective perturbation is optimized to induce misextraction of appearance features from the image and create incoherence among the generated video frames. Our extensive evaluation across 8 animation methods and 4 datasets demonstrates the superiority of Dormant over 6 baseline protection methods, leading to misaligned identities, visual distortions, noticeable artifacts, and inconsistent frames in the generated videos. Moreover, Dormant shows effectiveness on 6 real-world commercial services, even with fully black-box access.
Warning: This paper contains unfiltered images generated by LDMs that may be disturbing to some readers.
@@ 1 Introduction
While pose-driven human image animation methods have revolutionized video generation, they also significantly lower the barriers to creating deceptive and malicious human videos, raising serious concerns about unauthorized image usage. With just a single image of the victim, attackers can generate countless human videos, depicting the individual performing any action dictated by the pose input. For example, attackers can create dancing videos of celebrities to post on video platforms for commercial gains, fabricate fake videos of politicians to incite social harm, or produce NSFW videos that disrupt people’s normal lives [51]. These manipulated videos are easy and low-cost to generate, yet they severely violate the victim’s portrait and privacy rights and potentially cause considerable harm to the unsuspecting individual.
@@
Despite the severity of these hazards, there is a clear lack of research specifically aimed at preventing the misuse of pose-driven human image animation. To our knowledge, the most relevant concurrent work is VGMShield [50], which is also the only protection method designed to counteract malicious image-to-video generation. VGMShield applies protective perturbation to the image, leading to low-quality video generation by producing incorrect or bizarre frames. The protective perturbation is optimized by targeting the encoding process of Stable Video Diffusion (SVD) [7], deceiving both the image and video encoders into misinterpreting the input image. However, this deviation in the embedding space is not sufficient to effectively disrupt the appearance features contained in images, thus failing to adequately protect human portraits. Furthermore, VGMShield demonstrates limited effectiveness and poor transferability when defending against non-SVD-based video generation methods, including various pose-driven human image animation techniques.
@@
While few studies focus specifically on the misuse of image-to-video generation, several protection methods exist to defend against text/image-to-image generation [41, 44, 60, 61, 65, 73, 75, 77, 84]. These methods typically optimize protective perturbations by targeting the fine-tuning process of customization techniques [57], the encoding process of the Variational AutoEncoder [18], or the reverse process of the denoising UNet [27]. However, current protections against image generation are inadequate for defending against pose-driven human image animation, as they fail to effectively deceive the advanced feature extractors in animation methods to disrupt appearance features, and break the temporal consistency across frames that is essential in video generation.
@@
In this paper, we propose Dormant (Defending against Pose-driven Human Image Animation), to address this notable gap in effective protection methods that prevent unauthorized image usage in pose-driven human image animation, thereby safeguarding portrait and privacy rights. As illustrated in Figure 1, Dormant applies protective perturbation to the reference image, resulting in the protected image that appears visually similar to the original but leads to poor-quality video generation when used, causing issues including mismatched appearances, distorted visuals, noticeable artifacts and incoherent frames. We assume black-box access to the animation methods and models that the potential attacker might use, and employ a transfer-based strategy to optimize the protective perturbation according to our proposed objective function, utilizing publicly available pre-trained models as surrogates. The optimization objective includes 1) feature misextraction, which induces deviations in the latent representation and comprehensively disrupts the semantic and fine-grained appearance features of the reference image; 2) frame incoherence, which targets the appearance alignment and temporal consistency among the generated video frames. We further adopt Learned Perceptual Image Patch Similarity (LPIPS) [79], Expectation over Transformation (EoT) [5], and Momentum [17] to enhance the imperceptibility, robustness, and transferability of the protective perturbation, respectively.
@@
$\bullet$ We address the challenge of effectively preventing unauthorized image usage in pose-driven human image animation, safeguarding individual’s rights to portrait and privacy.
$\bullet$ We design novel objective functions to optimize the protective perturbation, inducing misextraction of appearance features and incoherence among generated video frames.
$\bullet$ Extensive experiments conducted on 8 pose-driven human image animation methods, 4 image-to-image methods, 4 image-to-video methods, and 6 commercial services highlight the effectiveness and transferability of Dormant [^1] Our code is available at https://github.com/Manu21JC/Dormant.
@@ 2.1 Latent Diffusion Model
Once trained, given $\tilde{z}_{T}$ sampled from the Gaussian distribution, the denoising UNet $\epsilon_{\theta}$ can be utilized to predict $\epsilon_{T}$ and computed $\tilde{z}_{T-1}$ according to Eq. (3). Subsequently, $\tilde{z}_{T-2},\tilde{z}_{T-3},\dots,\tilde{z}_{1},\tilde{z}_{0}$ are progressively calculated, and finally $\tilde{z}_{0}$ is decoded by $\mathcal{D}$ to generate the image $\tilde{x}=\mathcal{D}(\tilde{z}_{0})$.
@@ 2.2 LDM for Human Image Animation
Pose-driven human image animation [39, 72, 31, 86, 67, 69] is an image-to-video task aimed at generating a human video by animating a static human image based on a given pose sequence. As shown in Figure 1, the created video needs to accurately align the appearance details from the reference image while following the motion guidance from the pose sequence, which can be produced by DWPose [74], DensePose [23], or OpenPose [11]. Recently, LDM-based animation methods have demonstrated superior effectiveness in addressing challenges in appearance alignment and pose guidance, enabling the generation of realistic and coherent videos.
Using LDM as the backbone, Stable Diffusion (SD) has achieved remarkable success in text/image-to-image generation tasks [55, 78]. Pre-trained SD models effectively capture high-quality content priors. However, their network architecture is inherently designed for image generation, lacking the capability to handle the temporal dimension required for video generation. To extend foundational text/image-to-image models for image animation, many works [72, 31, 86, 15, 68] have incorporated temporal layers [24] into the denoising UNet for temporal modeling. Specifically, the feature map $X\in\mathbb{R}^{b\times f\times h\times w\times c}$ is reshaped to $X\in\mathbb{R}^{(b\times h\times w)\times f\times c}$, after which temporal attention is performed to capture the temporal dependencies among frames. Recently, some studies [52, 67, 82] have also adopted Stable Video Diffusion [7] as the backbone for image animation. SVD is an open-source image-to-video LDM trained on a large-scale video dataset, its strong motion-aware prior knowledge facilitates maintaining temporal consistency in image animation.
@@ 2.3 Protection Methods against LDM
While few studies addressing the issue of image-to-video generation, prior research has proposed various protection methods [84, 65, 44, 61, 60, 75, 41, 73, 77] against LDM to prevent image misuse in text/image-to-image generation. In the context of image generation, there are two primary LDM-based image misuse scenarios: concept customization and image manipulation. Existing protection methods mainly utilize Projected Gradient Descent (PGD) [45] to generate adversarial examples [21, 13] for LDM, thereby safeguarding against malicious image generation in both misuse scenarios. The defender aims to introduce protective perturbation $\delta$ to the image $x$, such that the protected image $x_{p}=x+\delta$ can deceive the LDM, resulting in poor-quality generation.
@@
There are several differences between concept customization, image manipulation, and pose-driven human image animation. First and foremost, concept customization and image manipulation are image generation tasks, whereas human image animation falls under the category of image-to-video generation. Second, concept customization requires several images (typically 3-5), while the other two techniques operate on a single image. Third, concept customization involves fine-tuning the LDM to memorize and generate a new concept. In contrast, image manipulation and human image animation leverage frozen pre-trained LDM directly for image or video generation without additional learning. Fourth, concept customization uses the images as training data for fine-tuning, image manipulation directly edits the given image itself, while human image animation extracts appearance features from the reference image to guide video generation, rather than using the image directly.
@@
While many protection methods exist to prevent unauthorized image usage in text/image-to-image generation, there is a notable gap in defending against image-to-video generation. To our knowledge, the only concurrent work addressing this issue is VGMShield [50], which aims to prevent malicious image-to-video generation using SVD [7] by generating adversarial examples. More precisely, it attacks the encoding process of SVD, aiming at deceiving the image encoder and video encoder into misinterpreting the image. However, this deviation specifically in the embedding space of SVD is insufficient to effectively disrupt the appearance features in human images and shows poor effectiveness and transferability when defending against non-SVD-based video generation methods, including various pose-driven human image animation techniques. Overall, the ineffectiveness of existing protections against image and video generation highlights the need for novel protection methods specifically designed to prevent malicious pose-driven human image animation.
@@ 3 Methodology
Figure 2: Overview of Dormant. Here, we present the four components of our proposed objective function $\mathcal{L}_{\textsc{Dormant}}$, which includes $\mathcal{L}_{vae}$ and $\mathcal{L}_{feature}$ for feature misextraction, $\mathcal{L}_{frame}$ for frame incoherence, and $\mathcal{L}_{lpips}$ for visual similarity.
Dormant prevents image misuse in pose-driven human image animation by adding protective perturbation $\delta$ to the human image $x$, resulting in the protected image $x_{p}=x+\delta$. $x_{p}$ is visually similar to $x$ but would lead to poor-quality video generation if used, causing issues such as mismatched appearances, visual distortions, and frame inconsistencies. We assume a fully black-box setting for the animation methods and models that potential attackers might use and employ a transfer-based strategy to optimize $\delta$ according to our proposed objective function $\mathcal{L}_{\textsc{Dormant}}$, relying on some surrogate pre-trained models (e.g., the publicly available CLIP model [53]) to which the defender has white-box access. As shown in Figure 2, $\mathcal{L}_{\textsc{Dormant}}$ contains: 1) $\mathcal{L}_{vae}$ and $\mathcal{L}_{feature}$, which aim to induce misextraction of appearance features from the reference image; 2) $\mathcal{L}_{frame}$, which is designed to cause incoherence among the generated video frames. Additionally, we adopt LPIPS [79], EoT [5], and Momentum [17], to further enhance the imperceptibility, robustness, and transferability of $\delta$, respectively. The protective perturbation is optimized using PGD [45], which iteratively updates $\delta$ with a step size $\gamma$ under an $L_{\infty}$ norm constraint $\eta$ as follows:
@@ 3.1 Threat Model
Attacker. The attacker aims to generate human videos using pose sequences based on the image obtained without the owner’s consent. These fabricated videos could be misused for commercial or political purposes, violating the victim’s rights to portrait and privacy. We assume the attacker maliciously acquires a single human image of the victim from the web or social media. Notably, even one image is enough to generate an unlimited number of videos. The attacker also has access to various LDMs for pose-driven human image animation, and significant computational power to run them.
Defender. We assume that the defender intends to apply protective perturbation to a human image before posting it online or on social media, to prevent potential unauthorized human video generation and thereby safeguard portrait and privacy rights. The protected image appears visually similar to the original image to human perception. However, any videos generated using this protected image would display mismatches in identity appearance and degraded quality, such as distorted backgrounds, visible artifacts, and incoherent frames, etc. To optimize the protective perturbation, the defender has access to some publicly available feature extractors and LDMs for pose-driven human image animation, but is unaware of the specific animation methods and models that the attacker might use. The defender could be either the image owner or a trustworthy third party with sufficient computing resources to execute the protection method effectively.
@@ 3.2 Feature Misextraction
In addition to using CLIP image encoder $\mathcal{C}$ to extract semantic features and optimize the protective perturbation $\delta$ according to Eq. (8), we also incorporate ReferenceNet $\mathcal{R}$ into the optimization process to enhance fine-grained feature extraction. To ensure that the protected image $x_{p}$ exhibits distinct detailed features from the perspective of ReferenceNet, we maximize the distance between the corresponding extracted features $\mathcal{R}(\mathcal{E}(x+\delta))$ and $\mathcal{R}(\mathcal{E}(x))$. Moreover, to further improve the transferability of the protective perturbation, we optimize $\delta$ using an ensemble [17, 42] of three different pre-trained ReferenceNets [72, 31, 15]. The objective $\mathcal{L}_{feature}$ for feature misextraction using the CLIP image encoder and ReferenceNets is thus defined as follows:
@@ 3.3 Frame Incoherence
While feature misextraction mainly focuses on disrupting the alignment of reference appearance, it is also crucial to consider the requirement of maintaining temporal consistency across frames in video generation when designing the optimization objective. Based on this, we further propose $\mathcal{L}_{frame}$, which attacks the video generation process of the denoising UNet to induce incoherence among the generated video frames, thus enhancing the protective effect. This requires the defender to provide the reference image as well as the pose sequence to simulate human image animation using LDM, just as the attacker does. However, it is hard for the defender to know in advance the exact pose sequence the attacker will use to generate the video. Therefore, we directly extract the corresponding pose from the reference image using DWPose [74] and repeat it $F$ times to serve as the pose guidance. In this case, each frame of the generated video should be identical to the others and consistent with the reference image.
@@
We set $F=5$ by default. As mentioned in Section 2.1, given the total inference steps $T$, we can sample $\tilde{z}_{T}$ from the Gaussian distribution and compute $\tilde{z}_{0}$ progressively from timestep $T$ to timestep 00 according to Eq. (3). However, this would result in substantial GPU memory usage and increased time costs. To improve efficiency, we directly estimate $\tilde{z}_{0}$ using the reparameterization technique outlined in Eq. (2), rather than computing it step by step. Specifically, given a timestep $t$, $\tilde{z}_{0}$ is estimated by $\tilde{z}_{0}=(\tilde{z}_{t}-\sqrt{1-\overline{\alpha}_{t}}\epsilon_{t})/\sqrt{\overline{\alpha}_{t}}$, where $\tilde{z}_{t}$ is sampled from the Gaussian distribution, and $\epsilon_{t}$ is predicted by a pre-trained denoising UNet [^3] https://github.com/MooreThreads/Moore-AnimateAnyone. In each PGD iteration, we randomly sample one timestep $t$ from the last ten timesteps of the total inference steps $T$ as larger $t$ makes better generation, and then estimate the corresponding $\tilde{z}_{0}$.

### v2

@@ Abstract
Pose-driven human image animation has achieved tremendous progress, enabling the generation of vivid and realistic human videos from just one single photo. However, it conversely exacerbates the risk of image misuse, as attackers may use one available image to create videos involving politics, violence, and other illegal content. To counter this threat, we propose Dormant, a novel protection approach tailored to defend against pose-driven human image animation techniques. Dormant applies protective perturbation to one human image, preserving the visual similarity to the original but resulting in poor-quality video generation. The protective perturbation is optimized to induce misextraction of appearance features from the image and create incoherence among the generated video frames. Our extensive evaluation across 8 animation methods and 4 datasets demonstrates the superiority of Dormant over 6 baseline protection methods, leading to misaligned identities, visual distortions, noticeable artifacts, and inconsistent frames in the generated videos. Moreover, Dormant shows effectiveness on 6 real-world commercial services, even with fully black-box access.
Warning: This paper contains unfiltered images generated by diffusion models that may be disturbing to some readers.
@@ 1 Introduction
While pose-driven human image animation methods have revolutionized video generation, they also significantly lower the barriers to creating deceptive and malicious human videos, raising serious concerns about unauthorized image usage [60, 49, 16]. With just a single image of the victim, attackers can generate countless human videos, depicting the individual performing any action dictated by the pose input. For example, attackers can create dancing videos of celebrities to post on video platforms for commercial gains, fabricate fake videos of politicians to incite social harm, or produce Not Safe for Work (NSFW) videos that disrupt people’s normal lives. These manipulated videos are easy and low-cost to generate, yet they severely violate portrait and privacy rights and potentially cause considerable harm to the unsuspecting individual.
@@
Despite the severity of these hazards, there is a clear lack of research specifically aimed at preventing the misuse of pose-driven human image animation. To our knowledge, the most relevant concurrent work is VGMShield [58], which is also the only protection method designed to counteract malicious image-to-video generation. VGMShield applies protective perturbation to the image, leading to low-quality video generation by producing incorrect or bizarre frames. The protective perturbation is optimized by targeting the encoding process of Stable Video Diffusion (SVD) [7], deceiving both the image and video encoders into misinterpreting the input image. However, this deviation in the embedding space is not sufficient to effectively disrupt the appearance features contained in images, thus failing to adequately protect human portraits. Furthermore, VGMShield demonstrates limited effectiveness and transferability when defending against non-SVD-based video generation methods, including various pose-driven human image animation techniques.
@@
While few studies focus specifically on the misuse of image-to-video generation, several protection methods exist to defend against text/image-to-image generation [48, 52, 70, 71, 76, 84, 86, 89, 96]. These methods typically optimize protective perturbations by targeting the fine-tuning process of customization techniques [67], the encoding process of the Variational AutoEncoder [20], or the reverse process of the denoising UNet [30]. However, current protections against image generation are inadequate for defending against pose-driven human image animation, as they fail to effectively deceive the advanced feature extractors in animation methods to disrupt appearance features and break the temporal consistency across frames that is essential in video generation.
@@
In this paper, we propose Dormant (Defending against Pose-driven Human Image Animation), to address this notable gap in effective protection methods that prevent unauthorized image usage in pose-driven human image animation, thereby safeguarding portrait and privacy rights. As illustrated in Figure 1, Dormant applies protective perturbation to the reference image, resulting in the protected image that appears visually similar to the original but leads to poor-quality video generation when used, causing issues including mismatched appearances, distorted visuals, noticeable artifacts, and incoherent frames. We assume that black-box access to the animation methods and models that the potential attacker might use, and employ a transfer-based strategy to optimize the protective perturbation according to our proposed objective function, utilizing publicly available pre-trained models as surrogates. The optimization objective includes: 1) feature misextraction, which induces deviations in the latent representation and comprehensively disrupts the semantic and fine-grained appearance features of the reference image; and 2) frame incoherence, which targets the appearance alignment and temporal consistency among the generated video frames. We further adopt Learned Perceptual Image Patch Similarity (LPIPS) [91], Expectation over Transformation (EoT) [5], and Momentum [19] to enhance the imperceptibility, robustness, and transferability of the protective perturbation, respectively.
@@
  • We address the challenge of effectively preventing unauthorized image usage in pose-driven human image animation, safeguarding individuals’ rights to portrait and privacy.
  • We design novel objective functions to optimize the protective perturbation, inducing misextraction of appearance features and incoherence among generated video frames.
  • Extensive experiments conducted on 8 pose-driven human image animation methods, 4 image-to-image methods, 4 image-to-video methods, and 6 commercial services highlight the effectiveness and transferability of Dormant.
@@ 2.1 Latent Diffusion Model
Once trained, given $\tilde{z}_{T}$ sampled from the Gaussian distribution, the denoising UNet $\epsilon_{\theta}$ can be utilized to predict $\epsilon_{T}$ and compute $\tilde{z}_{T-1}$ according to Eq. (3). Subsequently, $\tilde{z}_{T-2},\tilde{z}_{T-3},\dots,\tilde{z}_{1},\tilde{z}_{0}$ are progressively calculated, and finally $\tilde{z}_{0}$ is decoded by $\mathcal{D}$ to generate the image $\tilde{x}=\mathcal{D}(\tilde{z}_{0})$.
@@ 2.2 LDM for Human Image Animation
Pose-driven human image animation [43, 83, 34, 98, 79, 77] is an image-to-video task aimed at generating a human video by animating a static human image based on a user-defined pose sequence. As shown in Figure 1, the created video needs to accurately align the appearance details from the reference image while following the motion guidance from the pose sequence, which can be produced by DWPose [85] or DensePose [26]. Recently, LDM-based animation methods have demonstrated superior effectiveness in addressing challenges in appearance alignment and pose guidance, enabling the generation of realistic and coherent videos.
Using LDM as the backbone, Stable Diffusion (SD) has achieved remarkable success in text/image-to-image generation tasks [65, 90]. Pre-trained SD models effectively capture high-quality content priors. However, their network architecture is inherently designed for image generation, lacking the capability to handle the temporal dimension required for video generation. To extend foundational text/image-to-image models for image animation, many works [83, 34, 98, 14, 78] have incorporated temporal layers [27] into the denoising UNet for temporal modeling. Specifically, the feature map $X\in\mathbb{R}^{b\times f\times h\times w\times c}$ is reshaped to $X\in\mathbb{R}^{(b\times h\times w)\times f\times c}$, after which temporal attention is performed to capture the temporal dependencies among frames. Recently, some studies [61, 94, 77] have also adopted Stable Video Diffusion [7] as the backbone for image animation. SVD is an open-source image-to-video LDM trained on a large-scale video dataset; its strong motion-aware prior knowledge facilitates maintaining temporal consistency in image animation.
@@ 2.3 Protection Methods against LDM
While few studies addressing the issue of unauthorized image-to-video generation, prior research has proposed various protection methods [96, 76, 52, 71, 70, 86, 48, 84, 89] against LDM to prevent image misuse in text/image-to-image generation. In the context of image generation, there are two primary LDM-based image misuse scenarios: concept customization and image manipulation. Existing protection methods mainly utilize Projected Gradient Descent (PGD) [53] to generate adversarial examples [23, 12] for LDM, thereby safeguarding against malicious image generation in both misuse scenarios. The defender aims to introduce protective perturbation $\delta$ to the image $x$, such that the protected image $x_{p}=x+\delta$ can deceive the LDM, resulting in poor-quality generation.
@@
There are several differences between concept customization, image manipulation, and pose-driven human image animation. First, concept customization and image manipulation are image generation tasks, whereas human image animation is categorized as image-to-video generation. Second, concept customization requires multiple images (typically 3-5), while the other two techniques operate on a single image. Third, concept customization involves fine-tuning the LDM to memorize and generate a new concept. In contrast, image manipulation and human image animation leverage frozen pre-trained LDM directly for image or video generation without additional learning. Fourth, concept customization uses images as training data for fine-tuning, image manipulation directly edits the given image itself, while human image animation extracts appearance features from the reference image to guide video generation, rather than using the image directly.
@@
While many protection methods exist to prevent unauthorized image usage in text/image-to-image generation, there is a notable gap in defending against image-to-video generation. To our knowledge, the only concurrent work addressing this issue is VGMShield [58], which aims to prevent malicious image-to-video generation using SVD [7] by generating adversarial examples. More precisely, it attacks the encoding process of SVD, aiming at deceiving the image encoder and video encoder into misinterpreting the image. However, this deviation specifically in the embedding space of SVD is insufficient to effectively disrupt the appearance features in human images and shows limited effectiveness and transferability when defending against non-SVD-based video generation methods, including various pose-driven human image animation techniques. Overall, the ineffectiveness of existing protections against image and video generation highlights the need for novel protection methods specifically designed to prevent malicious pose-driven human image animation.
@@ 3 Methodology
Figure 2: Overview of Dormant. Here, we present the four components of our proposed objective function $\mathcal{L}_{\textsc{Dormant}}$, which includes: $\mathcal{L}_{vae}$ and $\mathcal{L}_{feature}$ for feature misextraction, $\mathcal{L}_{frame}$ for frame incoherence, and $\mathcal{L}_{lpips}$ for visual similarity.
Dormant prevents image misuse in pose-driven human image animation by adding protective perturbation $\delta$ to the human image $x$, resulting in the protected image $x_{p}=x+\delta$. $x_{p}$ is visually similar to $x$ but would lead to poor-quality video generation if used, causing issues such as mismatched appearances, visual distortions, and frame inconsistencies. We assume that a fully black-box setting for the animation methods and models that potential attackers might use and employ a transfer-based strategy to optimize $\delta$ according to our proposed objective function $\mathcal{L}_{\textsc{Dormant}}$, relying on some surrogate pre-trained models (e.g., the publicly available CLIP model [62]) to which the defender has white-box access. As shown in Figure 2, $\mathcal{L}_{\textsc{Dormant}}$ contains: 1) $\mathcal{L}_{vae}$ and $\mathcal{L}_{feature}$, which aim to induce misextraction of appearance features from the reference image; and 2) $\mathcal{L}_{frame}$, which is designed to cause incoherence among the generated video frames. Additionally, we adopt LPIPS [91], EoT [5], and Momentum [19], to further enhance the imperceptibility, robustness, and transferability of $\delta$, respectively. The protective perturbation is optimized using PGD [53], which iteratively updates $\delta$ with a step size $\gamma$ under an $L_{\infty}$ norm constraint $\eta$ as follows:
@@ 3.1 Threat Model
Attacker. The attacker aims to generate human videos using pose sequences based on images obtained without the owner’s consent. These fabricated videos could be misused for commercial or political purposes, violating the victim’s rights to portrait and privacy. We assume that the attacker maliciously acquires multiple images of the victim from the web or social media. The attacker is aware of potential protections and can apply various countermeasures to these images to craft a suitable reference image for video generation. The attacker also has access to various LDMs for pose-driven human image animation, and significant computational power to run them.
Defender. We assume that the defender intends to apply protective perturbation to a human image before posting it online or on social media, to prevent potential unauthorized human video generation and thereby safeguard portrait and privacy rights. The protected image appears visually similar to the original image to human perception. However, any videos generated using this protected image would display mismatches in identity appearance and degraded quality, such as distorted backgrounds, visible artifacts, and incoherent frames, etc. To optimize the protective perturbation, the defender has access to some publicly available feature extractors and LDMs for pose-driven human image animation. However, the defender is unaware of, and therefore unable to approximate, the specific animation methods and models that the attacker might use. The defender could be either the image owner or a trustworthy third party with sufficient computing resources to execute the protection method effectively.
@@ 3.2 Feature Misextraction
In addition to using CLIP image encoder $\mathcal{C}$ to extract semantic features and optimize the protective perturbation $\delta$ according to Eq. (8), we also incorporate ReferenceNet $\mathcal{R}$ into the optimization process to enhance fine-grained feature extraction. To ensure that the protected image $x_{p}$ exhibits distinct detailed features from the perspective of ReferenceNet, we maximize the distance between the corresponding extracted features $\mathcal{R}(\mathcal{E}(x+\delta))$ and $\mathcal{R}(\mathcal{E}(x))$. Moreover, to further improve the transferability of the protective perturbation, we optimize $\delta$ using an ensemble [19, 50] of $K=3$ different pre-trained ReferenceNets [83, 34, 14]. The objective $\mathcal{L}_{feature}$ for feature misextraction using the CLIP image encoder and ReferenceNets is thus defined as follows:
@@ 3.3 Frame Incoherence
While feature misextraction mainly focuses on disrupting the alignment of reference appearance, it is also crucial to consider the requirement of maintaining temporal consistency across frames in video generation when designing the optimization objective. Based on this, we further propose $\mathcal{L}_{frame}$, which attacks the video generation process of the denoising UNet to induce incoherence among the generated video frames, thus enhancing the protective effect. This requires the defender to provide both the reference image and the pose sequence to simulate the attacker’s animation process. However, predicting the exact pose sequences the attacker might use is challenging. To this end, we directly extract the corresponding pose from the reference image and repeat it $F$ times as the pose sequence to guide video generation.
@@
We set $F=5$ by default. As mentioned in Section 2.1, given the total inference steps $T$, we can sample $\tilde{z}_{T}$ from the Gaussian distribution and compute $\tilde{z}_{0}$ progressively from timestep $T$ to timestep 00 according to Eq. (3). However, this would result in substantial GPU memory usage and increased time costs. To improve efficiency, we directly estimate $\tilde{z}_{0}$ using the reparameterization technique outlined in Eq. (2), rather than computing it step by step. Specifically, given a timestep $t$, $\tilde{z}_{0}$ is estimated by $\tilde{z}_{0}=(\tilde{z}_{t}-\sqrt{1-\overline{\alpha}_{t}}\epsilon_{t})/\sqrt{\overline{\alpha}_{t}}$, where $\overline{\alpha}_{t}=\prod^{t}_{i=1}(1-\beta_{i}$), $\{\beta_{t}\in(0,1)\}^{T}_{t=1}$ is the variance schedule, $\epsilon_{t}$ is predicted by a pre-trained denoising UNet [^2] https://github.com/MooreThreads/Moore-AnimateAnyone, and $\tilde{z}_{t}$ is sampled from the Gaussian distribution. In each PGD iteration, we randomly sample one timestep $t$ from the last 10 timesteps of $T$ and estimate the corresponding $\tilde{z}_{0}$.